OpenStudy (anonymous):

convert 144 ft; inches

OpenStudy (anonymous):

144 feet = 1728 inches

jimthompson5910 (jim_thompson5910):

1 ft = 12 inches 144*1 ft = 144*12 inches 144 ft = 1728 inches